OpenStudy (anonymous):

x – 9 = –15 A. –6 B. –24 C. 24 D. 6

OpenStudy (love_333):

A! :)

OpenStudy (netlopes1):

\[x-9=-15\rightarrow x=-15+9\rightarrow x=-6\]

OpenStudy (netlopes1):

But do you know why? or not...